Compute fxy and fyx for the function f (x,y)=e^x+y sinx +9x^2+2y at (1,2)
1
Expert's answer
2020-03-16T13:38:30-0400

"f (x,y)=e^x+y sinx +9x^2+2y"

"f_{xy}=(f_x)_y= \frac{\partial }{\partial y}(\frac{\partial f}{\partial x})"

"=\frac{\partial }{\partial y}(e^x+ycosx+18x)"

"=cosx"

Thus, "f_{xy}(1,2)=\cos(1)"


"f_{yx}=(f_y)_x= \frac{\partial }{\partial x}(\frac{\partial f}{\partial y})"

"=\frac{\partial }{\partial x}(sinx+2)"

"=cosx"

Thus, "f_{yx}(1,2)=cos(1)"